### v1

@@ 8 Conclusion
In this study, we introduce a multimodal data retrieval platform MQRLD. This platform builds on a data lake which offers transparent data storage and integrates a multimodal open API for rich hybrid queries. In addition to these capabilities, its multimodal data representation strategy transforms data into effective feature vectors through feature embedding, measurement, and enhancement, aiding accurate data retrieval. To further improve query efficiency, high-dimensional learned indexes are built using divisive hierarchical clustering and cluster tree construction, with the query-aware mechanism optimizing the index structure. Overall, MQRLD not only supports transparent data storage and rich hybrid queries, but also ensures effective and efficient retrieval.
Our study provides a broad scope for future work in multimodal data management. Additionally, the optimization strategies we propose for the multimodal data retrieval platform also have versatile applications. Experiments have demonstrated that our proposed multimodal data representation strategy and high-dimensional learned index effectively improve retrieval accuracy and efficiency on data lakes. Despite the limitations of existing platforms in multimodal data retrieval, their inherent features offer potential in specific scenarios, as discussed in Section 1. Therefore, we consider applying our multimodal data representation strategy and high-dimensional learned index to other platforms in the future to enhance their multimodal data retrieval performance.
@@ Acknowledgements
The work is funded by the National Natural Science Foundation of China under Grant 62076027.

### v2

@@ 8 Conclusion
In this study, we introduce a multimodal data retrieval platform MQRLD. This platform builds on a data lake that offers transparent data storage and integrates a multimodal open API for rich hybrid queries. In addition to these capabilities, its multimodal data representation strategy transforms data into effective feature vectors through feature embedding, measurement, and enhancement, aiding accurate data retrieval. To further improve query efficiency, high-dimensional learned indexes are built using divisive hierarchical clustering and cluster tree construction, with the query-aware mechanism optimizing the index structure. Overall, MQRLD not only supports transparent data storage and rich hybrid queries but also ensures effective and efficient retrieval.
Although our proposed MQRLD addresses the functionalities of transparent storage and rich hybrid queries, and enhances query performance, as mentioned in Section 1, it still faces additional challenges: (1) The MQRLD platform mainly consists of two modules: feature representation and high-dimensional learned index. Experiments have shown that these modules can effectively improve the retrieval performance of data lakes. However, the scalability of them on other platforms still needs further discussion. (2) Despite the large accumulation of multimodal data in the form of historical records, the development of Internet of Things (IoT) technologies supports multimodal data in a streaming format. Furthermore, user query behaviors often exhibit varying preferences over time. Therefore, to maintain the platform retrieval performance, whether it can support dynamic index reconstruction when facing frequent data appending and different query workloads remains a challenge. (3) Large language models (LLMs) have demonstrated their high scalability in different fields. It remains to be explored whether MQRLD can leverage LLMs to achieve more efficient and high-quality feature representation. (4) Although multimodal data are primarily accumulated in batch form, it is reasonable to assume that more recent data are likely to exhibit more similar characteristics. Additionally, query workloads within a specific time period are likely to exhibit similar preferences. Therefore, it is worth exploring whether temporal information can be utilized by the platform to enhance feature representation and index construction.
We can further optimize our platform in the future, including: (1) We consider applying our multimodal data feature representation strategy and high-dimensional learned index to other platforms to enhance their multimodal data retrieval performance while preserving their inherent characteristics as discussed in Section 1. (2) Existing researches typically maintain retrieval performance under frequent data appending and different query workloads by monitoring query results to determine whether the index needs to be reconstructed (Ding et al., 2021; Xie et al., 2023). So how to determine the threshold for triggering index reconstruction is an opportunity worth exploring. Moreover, to avoid the expensive cost of reconstructing the entire index, it is worth considering whether partial index optimizations can help balance the trade-off among online data appending, query workloads, and index reconstruction. (3) LLMs demonstrate significant potential in representation learning. Currently, the feature embedding models pool in our feature representation includes the popular CLIP-based models (Radford et al., 2021; Luo et al., 2021; Ma et al., 2022). However, these models do not support a unified training representation across various modalities. Exploring the possibility of leveraging LLMs to offer a unified embedding strategy of different modalities presents an interesting opportunity for further investigation. (4) Liu et al. (2023b) shows that incorporating temporal information into multimodal graph learning can effectively aid multimodal data classification tasks, which may help with feature representation (aims to gather similar multimodal data). Additionally, Liu et al. (2023c) indicates that temporal information can enhance the prediction of interactions, supporting the implementation of online query workloads on our platform (allowing the platform to dynamically adjust the index in advance based on predicted query workload (Pavlo et al., 2021)). Therefore, leveraging temporal information to optimize our platform functionalities and performance presents a valuable opportunity.
@@ Acknowledgements
The work is funded by the National Natural Science Foundation of China under Grant 42371480.